OpenStudy (anonymous):

The function "f" is differentiable and (4f(t)+3t)dt over interval [0,x] = sin(x). Determine f'(π/6).

OpenStudy (irishboy123):

using \[\int_{0}^{x} \ g(t) \ dt = G(x) - G(0)\] meaning \[\frac{d}{dx} \int_{0}^{x} \ g(t) \ dt = \frac{d}{dx} [G(x) - G(0)] = g(x)\] we have: \[\int_{0}^{x} \ 4f(t) + 3t \ dt = sin(x) \] meaning \[\frac{d}{dx} \int_{0}^{x} \ 4f(t) + 3t \ dt = \frac{d}{dx} sin(x) \] and \[4f(x) + 3x = cos x\] from that, find \( f'(x)\) and then \( f'(\pi/6)\)